Given the Earnings before interest and taxes(EBIT) is Rs 30000 Interest payment is Rs 10000

Dividend on the preference shares Rs9000

Taxes being 50% of the profit before taxes(PBT).

Number of outstanding equity shares 10000

  Weightage

For Numerical Answer

Assessment Parameter Weightage Understanding and usage 20%

of the formula

Procedure / Steps 50%

  20% Conclusion 20%

Correct Answer Interpretation

& 30%

What would be the earning per share(EPS) and degree of financial leverage (DFL), What would be the change in EPS, and DFL, if the EBIT increases to Rs50000 and Rs80000

Note- you have to calculate three EPS and three calculations for DFL



1
Expert's answer
2021-10-05T09:21:00-0400

Solution:

EPS = "\frac{Net\; income}{No.\; of \; outstanding \; equity \; shares}"

No. of outstanding equity shares = 10,000

Net income = EBIT – Interest – Taxes

EBT or PBT = EBIT – Interest = 30,000 – 10,000 = 20,000

Taxes = 50"\% \times" 20,000 = 10,000

Net income = 20,000 – 10,000 = 10,000

EPS = 10,000/10,000 = 1

 

DFL = EBIT/EBIT – Interest

= "\frac{EBIT}{EBIT - Interest} =\frac{30,000}{30,000 - 10,000} = \frac{30,000}{20,000} = 1.5"

 

When EBIT increases to 50,000:

EBT or PBT = EBIT – Interest = 50,000 – 10,000 = 40,000

Taxes = 50"\% \times" 40,000 = 20,000

Net income = 40,000 – 20,000 = 20,000

EPS = "\frac{20,000}{10,000} = 2"

 

DFL = "\frac{EBIT}{EBIT - Intrest} = \frac{50,000}{50,000 - 10,000} = \frac{50,000}{40,000} = 1.25"

 

When EBIT increases to 80,000:

EBT or PBT = EBIT – Interest = 80,000 – 10,000 = 70,000

Taxes = 50"\% \times" 70,000 = 35,000

Net income = 70,000 – 35,000 = 35,000

EPS = "\frac{35,000}{10,000} = 3.5"

 

DFL = "\frac{EBIT}{EBIT - Interest}" = "\frac{80,000}{80,000 - 10,000} = \frac{80,000}{70,000} = 1.14"